OpenStudy (anonymous):

Suppose the circumference of a circle is 14π. What is its radius?

OpenStudy (astrophysics):

\[\huge C = 2 \pi r\] You're given the circumference C is 14 pi, so solve for here \[\huge 14 \pi = 2 \pi r\]

OpenStudy (astrophysics):

Solve for r

rebeccaxhawaii (rebeccaxhawaii):

14/2=?